# Context-Aware Replanning with Pre-explored Semantic Map for Object Navigation

### v1

@@ Abstract
Pre-explored Semantic Maps, constructed through prior exploration using visual language models (VLMs), has proven effective as foundational elements for training-free robotic applications. However, existing approaches assume the map’s accuracy and do not provide effective mechanisms for revising decisions based on incorrect maps. To address this, we introduce Context-Aware Replanning (CARe), which estimates map uncertainty through confidence scores and multi-view consistency, enabling the agent to revise erroneous decisions stemming from inaccurate maps without requiring additional labels. We demonstrate the effectiveness of our proposed method by integrating it with two modern mapping backbones, VLMaps and OpenMask3D, and observe significant performance improvements in object navigation tasks. More details can be found on the project page: https://carmaps.github.io/supplements/.
Keywords: Object Navigation, Map Uncertainty Measurement, Replanning
@@ 1 Introduction
Object navigation, which involves localizing and navigating to an object in indoor environments, is a critical component of robotic applications. Conventional training-based methods necessitate extensive annotations, meticulous model design, and prolonged training periods to effectively align the control actions with visual perception. Advancements in visual language models (VLMs) have led to the development of modular approaches that separate perception from actions, utilizing pre-trained knowledge. Under this framework, visual perception can be independently learned without direct control, making exploration prior to task execution an effective strategy. Pre-explored Semantic Map [1, 2, 3, 4, 5], constructed through prior exploration and using visual language models (VLMs), has become a fundamental backbone for robotics tasks. By constructing the map during environmental exploration, Pre-explored Semantic Map equips the agent with pre-existing knowledge of the environment, facilitating training-free robotic tasks such as manipulation motion planning [6], interactive exploration [7], and zero-shot object navigation [5]. However, current approaches presume Pre-explored Semantic Map is always accurate and lacks effective mechanisms for revising decisions based on incorrect maps. This assumption is incorrect for real-world applications, as visual perception must contend with varied environments featuring different lighting conditions, textures, weather, and dynamic elements. Consequently, visual perception cannot be assumed to be perfect. In addition, evaluating the quality of the Pre-explored Semantic Map is also challenging, as there are no existing labels for comparison.
@@
Intuitively, the map-based agent plans its path by retrieving the goal from the map, where retrieval is facilitated by sorting confidence scores based on map-query matching. Therefore, should the initial attempt fail, the agent can proceed to another unvisited location with the highest confidence score. Nevertheless, this approach assumes the map is accurate. However, failure to retrieve the map with the highest confidence score may indicate a mismatch or error in the map-query alignment. Given the observation that the visual observation upon which the map is based fails to satisfy the query, we hypothesize that integrating Uncertainty measurement into the decision-making process can enhance the agent’s ability to cope with inaccuracies in the map. By prioritizing areas with the greatest uncertainty, the agent targets high-uncertainty points for maximum information gain. This improves performance because initial prediction failures often result from biases or errors in the map data. Replanning based solely on confidence scores can perpetuate these biases. High-uncertainty areas, however, are less influenced by existing biases, providing new, unbiased information that can correct or enhance the agent’s understanding. Thus, integrating uncertainty measurement reduces bias impact and increases the likelihood of successfully finding the goal. In addition, the same region can be observed from multiple viewpoints. If the VLM predictions from these multiple viewpoints are not aligned, the prediction is likely incorrect. Therefore, Multi-view Consistency can also serve as an unsupervised metric for agent replanning.
@@
Figure 1: When the initial map query fails, high-confidence regions also tend to fail due to biases in visual perception. (Query: Table, Grounding: Chair)
@@
To this end, we propose a novel Context-Aware Replanning (CARe) to revise the plan when the initial planning fails. Specifically, we first select $k$ candidate regions as a strategy to mitigate the limitations of imperfect perception systems. While the top-ranked region might not always be reliable due to potential inaccuracies in perception, selecting a broader set of top candidates ensures a more robust assessment. Afterward, we measure uncertainty and multi-view consistency for candidate regions. For uncertainty, we sort the regions by the entropy of VLM predictions and select the highest. For multi-view consistency, we measure the average pair-wise KL-divergence and select the lowest. Context-Aware Replanning is modular, portable to different backbones, and leverages the context of Pre-explored Semantic Map without requiring additional annotations. Consequently, Context-Aware Replanning enhances decision-making resilience and, while focused on replanning, holds potential for continuous map refinement in future studies.
To demonstrate the modularity and robustness of CARe, we conducted experiments using two Pre-explored Semantic Map backbones, VLMaps [5] and OpenMask3D [3], showcased their effectiveness in object navigation tasks. The results demonstrate that CARe consistently outperforms the strategy of selecting the highest-scored unvisited region. We tested two different strategies for measuring uncertainty: single-view confidence-based and multi-view consistency. Both approaches proved effective, validating our hypothesis that leveraging uncertainty improves the overall effectiveness of Pre-explored Semantic Map-based applications, with navigation serving as a prime example. Given its modularity and portability, we encourage future research to consider integrating CARe into a broader range of robotic applications.
The contribution of this work is summarized as follow:
@@
  • Our method consistently outperforms the strategy of selecting the highest-scored unvisited region using two different backbones, VLMaps [5] and OpenMask3D [3], demonstrating its effectiveness and robustness.
@@ 2 Related Work
Visual Language Models (VLMs) serve as the backbone of modern modular training-free robotic applications by aligning visual perception with language, as instructions are typically given in natural language. CLIP [8] achieves this alignment through contrastive learning, forming the foundation for various training-free robotic applications. For instance, OpenMask3D [3] and CLIP-Fields [2] utilize CLIP to construct maps. Additionally, several open-vocabulary visual perception tools, such as GLIP [9] for object detection and LSeg [10] for semantic segmentation, are crucial for robotic applications. VLMaps [5] uses LSeg to build its maps. While VLMs achieve significant performance on various computer vision tasks, robotic applications must handle diverse environments with varying lighting conditions, textures, weather, and object categories. Therefore, accounting for and replanning with errors in maps is crucial when leveraging maps constructed with VLMs.
@@ 3 Method
In this section, we first introduce the background of our method in 3.1. Next, we illustrate the motivating concept behind our method to address a significant issue in current approaches in Section 3.2. We then detail our approach, which primarily focuses on leveraging uncertainty measures to select from a set of high-confidence candidates, as described in Section 3.3. Within this section, we further explain the single and multi-view uncertainty measures that our method utilizes. Subsequently, we outline the strategies we tested to generate the set of high-confidence candidates in Section 3.4. Finally, we describe the maps used in our experiments and discuss how they were adapted to integrate with our method in Section 3.5.
@@
The conventional way to use navigation maps typically involves using a text query to retrieve the candidate with the highest matching score. Such a score is typically calculated with pretrained visual-language grounding models such as CLIP[8]. These grounding models have demonstrated impressive generalization abilities due to the large dataset used in the pertaining phase. This is probably why most of the works navigating with maps assume that the map is perfect. Nevertheless, we argue that despite the size of the training dataset, there might still be bias introduced by the model architecture or the dataset, which causes the user expectation of a query to diverge from the model’s belief. Furthermore, during the construction phase of the map, the model prediction might be interfered with by surrounding objects or some view-dependent bias, causing the extracted feature to be noisy.
@@ 3.3.1.1 Entropy:
Let $C=\{c_{1},c_{2},\ldots,c_{n}\}$ be the set of high-confidence candidates. For each candidate $c_{i}$, we compute the entropy $H(c_{i})$ of its classification distribution $P_{i}=\{p_{1},p_{2},\ldots,p_{k}\}$ as:
@@
where $p_{j}$ is the probability of candidate $c_{i}$ belonging to class $j$.
Finally, we select the candidate $c^{*}$ with the highest entropy:
@@
With this definition, a high entropy indicates that the classification distribution resembles a spread-out probability distribution, suggesting that the model considers several classes as plausible alternatives, thereby reflecting uncertainty.
This approach ensures that we are not only considering candidates with high confidence but also those with high uncertainty, thereby reducing the likelihood of biased retrievals. For the specific method of selecting high-confidence candidates, please refer to Section 3.4.
@@
When the feature of a candidate is extracted by a small number of views, the resulting feature can easily be interfered by nearby objects or view-dependent model bias. To mitigate this issue, we propose to encourage multi-view consistency when retrieving the candidate. Similar to the single-view case, we first identify a list of high-confidence candidates and then encourage the retrieval of low-multi-view inconsistency (uncertainty) candidates. We experimented two kinds of multi-view uncertainty measures: channel-average feature standard error and mean pairwise KL divergence on multi-view classification probabilities.
@@ 3.3.2.1 Channel-average Feature Standard Error:
where $\sigma_{k}$ is the standard deviation of the $k$-th feature across all views, $d$ is the dimensionality of the feature vector, and $m_{i}$ is the number of views for candidate $c_{i}$. And in the case of VLMaps, we incorporate the distance weighting [1, 5] information of each feature which is utilized when building the map, that is we replace the standard deviation and the sample size with the weighted version:
$SE^{w}(c_{i})=\frac{1}{d}\sum_{k=1}^{d}\frac{\sigma^{w}_{k}}{\sqrt{m^{eff}_{i}}}$
$m^{eff}_{i}=\frac{(\sum_{i=1}^{n}w_{i})^{2}}{\sum_{i=1}^{n}{w_{i}^{2}}}$
Finally, we choose the candidate $c^{*}$with the lowest standard error.
Since most maps utilize the mean feature from multiple views for retrieval, a low standard error on these multi-view features indicates that the mean feature is statistically more reliable. Therefore encouraging low standard error could potentially reduce noisy retrievals.
@@ 3.3.2.2 Mean Pairwise KL Divergence:
Finally, we choose the candidate $c^{*}$with the lowest mean pairwise KL divergence.
A low mean pairwise KL divergence indicates that the classification distribution is consistent across multiple views. Similar to the standard error, promoting consistency in the classification distribution could help reduce noisy retrievals.
@@ 3.4.2 Top-k Category:
OpenMask3D[11] takes in the scene point cloud and posed RGB images, generates class-agnostic 3D masks on the point cloud, and uses the posed RGB images where the object is visible to provide semantic features. The features are calculated with CLIP with cropped images of the corresponding object and thus enable retrieval later on. If the object is visible in multiple views, OpenMask3D takes the average of multi-view features.
We augmented OpenMask3D to not take the average of the multi-view features, but simply record features from all visible views for each object mask.
Let:
  • $P$ be the scene point cloud.
  • $I=\{I_{1},I_{2},\ldots,I_{n}\}$ be the set of posed RGB images.
  • $V_{i}$ be the set of views where object $m_{i}$ is visible.
$F_{ij}=\text{CLIP}(\text{crop}(I_{j},m_{i})),\quad\forall I_{j}\in V_{i}$
where $F_{ij}$ is the feature vector for object $m_{i}$ from view $I_{j}$.
$F_{i}=\{F_{ij}\mid I_{j}\in V_{i}\}$
@@
VLMaps takes the RGB-D image and its corresponding pose as input. It first generate the local point cloud which is then projected to the world coordinate frame and the map position with the information of camera poses. After that, the RGB image is fed into Visual Encoder of LSeg[10] to get its image feature, which will then be projected to the corresponding map position. Though we do not modify the process of building the map, we store some more data such as the standard error and KL divergence of each point on the map to decrease the size of the map and avoid repetitive computation in the evaluation process.
@@ 4 Experiments
Table 1: OM3D Object Retrieval Success Rates. The upper half of the table presents the baseline methods, as described in 4.1.2, while the lower half displays variants of our method. The ’Replan Strategy’ column indicates the replanning strategies used, as detailed in 3.3. The ’Selection Criteria’ column specifies the criteria employed to generate the high-confidence candidate set, as described in 3.4. The best-performing entry for each column, excluding the Oracle, is bolded to emphasize the best-performing method given $k$. The best-performing entry for each row is highlighted to showcase the performance of each method when a hyper-parameter search on $k$ is available. Finally, the best-performing baseline entry is italicized for easier comparison.
@@
4.3.1 Baselines.
  • No replan: This is the success rate without any replanning.
  • Max confidence (top2 score): This is a naive baseline where the point with the largest score is selected when the first trial fails, which matches the typical usage of navigation maps.
  • Random replan: This baseline randomly selects a candidate as the second attempt.
@@
Table 2 showcases the effectiveness of various replanning strategies compared to the ”Max confidence (highest score)” baseline. The ”Max entropy from top-k confidence” and ”Max entropy from top-k category” strategies both demonstrate significant improvements over the baseline, reducing bias by considering entropy.
The ”Min stderr” strategy along with ”confidence” selection criteira consistently outperforms all other methods, including the baseline, by focusing on minimizing the standard error in confidence predictions. This approach proves to be the most robust and effective, highlighting the importance of incorporating uncertainty and multi-view consistency in replanning for superior outcomes.
@@
Table 2: VLMaps Replanning Subgoal Success Rates. Similar to table 1, the baseline methods and ’Replan Strategy’ column are explained in 4.3.1 and 3.3, respectively. The ’Selection Criteria’ column describes the criteria for generating the high-confidence candidate set, as detailed in 3.4. The best-performing entry in each column is bolded and in each row is highlighted . The best-performing baseline entry is italicized for easier comparison.
4.4 Ablation Studies.
4.4.1 Different Choice on choosing Maximal/Minimal Uncertainty.
To further support our assumptions, we conducted an ablation study with the OpenMask3D setting by examining different directions of the uncertainty metric.
Single-view Entropy:
If a model is biased, selecting low-entropy candidates from it might further reinforce the bias and ultimately degrade performance. Figure 3 illustrates the performance of choosing minimum entropy targets from top-k confidence candidates. This approach consistently underperforms the random baseline.
Multi-view Consistency:
When scores from different views are inconsistent, it may indicate that the resulting scores are noisier and more likely to provide false information. Figure 3 depicts the performance of selecting the least consistent candidate (i.e., maximum mean pairwise KL divergence) from top-k confidence candidates. Similar to the single-view case, this variant, which discourages multi-view consistency, underperforms the baselines.
Figure 2: Entropy vs. K
Figure 3: KL Divergence vs. K

### v2

@@ Abstract
Pre-explored Semantic Maps, constructed through prior exploration using visual language models (VLMs), have proven effective as foundational elements for training-free robotic applications. However, existing approaches assume the map’s accuracy and do not provide effective mechanisms for revising decisions based on incorrect maps. To address this, we introduce Context-Aware Replanning (CARe), which estimates map uncertainty through confidence scores and multi-view consistency, enabling the agent to revise erroneous decisions stemming from inaccurate maps without requiring additional labels. We demonstrate the effectiveness of our proposed method by integrating it with two modern mapping backbones, VLMaps and OpenMask3D, and observe significant performance improvements in object navigation tasks. More details can be found on the project page: https://care-maps.github.io/.
Keywords: Object Navigation, Vision-Language Models, Uncertainty Measurement, Training-Free Replanning
@@ 1 Introduction
Navigating in indoor environments to locate and reach target objects is a fundamental capability for robotic applications. Conventional training-based object navigation methods necessitate extensive annotations, meticulous model design, and prolonged training periods to effectively align the control actions with visual perception. Advancements in visual language models (VLMs) have led to the development of modular approaches that separate perception from actions, utilizing pre-trained knowledge. Under this framework, visual perception can be independently learned without direct control, making exploration prior to task execution an effective strategy. Pre-explored Semantic Map [1, 2, 3, 4, 5], constructed through prior exploration and using visual language models (VLMs), has become a fundamental backbone for robotics tasks. By constructing the map during environmental exploration, Pre-explored Semantic Map equips the agent with pre-existing knowledge of the environment, facilitating training-free robotic tasks such as manipulation motion planning [6], interactive exploration [7], and zero-shot object navigation [5]. However, current approaches presume that Pre-explored Semantic Map is always accurate—unbiased and noise-free—and they lack effective mechanisms for revising decisions based on incorrect maps. This assumption is flawed for real-world applications, where visual perception must contend with diverse environments, including varying lighting conditions, textures, weather, and dynamic elements. Consequently, visual perception cannot be assumed to be perfect. In addition, evaluating the quality of the Pre-explored Semantic Map is also challenging, as there are no existing labels for comparison.
@@
Intuitively, the map-based agent plans its path by retrieving the goal from the map, where retrieval is facilitated by sorting confidence scores based on map-query matching. Therefore, should the initial attempt fail, the agent can proceed to another unvisited location with the highest confidence score. Nevertheless, this approach assumes the map is accurate. However, failure to retrieve the map with the highest confidence score may indicate a mismatch or error in the map-query alignment. Given the observation that the visual observation upon which the map is based fails to satisfy the query, we hypothesize that integrating Uncertainty measurement into the decision-making process can enhance the agent’s ability to cope with inaccuracies in the map. By prioritizing areas with the greatest uncertainty, the agent targets high-uncertainty points for maximum information gain. This improves performance because initial prediction failures often result from biases or errors in the map data. Replanning based solely on confidence scores can perpetuate these biases. High-uncertainty areas, however, are less influenced by existing biases. Therefore, integrating uncertainty measurement reduces bias impact and increases the likelihood of successfully finding the goal. In addition, the same region can be observed from multiple viewpoints. If the VLM predictions from these multiple viewpoints are not aligned, the prediction is likely incorrect. Therefore, Multi-view Consistency can also serve as an unsupervised metric for agent replanning.
@@
Figure 1: Motivation. When the initial map query fails, high-confidence regions also tend to fail due to biases in visual perception. (Query: Table, Grounding: Chair)
Figure 2: Method Overview.
@@
To this end, we propose a novel Context-Aware Replanning (CARe) to revise the plan when the initial planning fails. Specifically, we first select $k$ candidate regions as a strategy to mitigate the limitations of imperfect perception systems. While the top-ranked region might not always be reliable due to potential inaccuracies in perception, selecting a broader set of top candidates ensures a more robust assessment. Afterward, we measure uncertainty and multi-view consistency for candidate regions and select the best candidate based on the measured values. Our evaluation in two popular Pre-explored Semantic Map backbones, VLMaps [5] and OpenMask3D [3], showcased that our method consistently outperforms the strategy of selecting the highest-scored unvisited region.
The contribution of this work is summarized as follows:
@@
  • Our method consistently outperforms the strategy of selecting the highest-scored unvisited region using two different backbones [5, 3], demonstrating its effectiveness and robustness.
@@ 2 Related Work
Visual Language Models (VLMs) serve as the backbone of modern modular training-free robotic applications by aligning visual perception with language, enabling robots to understand and follow natural language instructions. CLIP [8] achieves this alignment through contrastive learning, forming the foundation for various training-free robotic applications. For instance, OpenMask3D [3] and CLIP-Fields [2] utilize CLIP to construct maps. Additionally, several open-vocabulary visual perception tools, such as GLIP [9] for object detection and LSeg [10] for semantic segmentation, are crucial for robotic applications. For instance, VLMaps [5] uses LSeg to build its maps. While VLMs achieve significant performance on various computer vision tasks, robotic applications must handle diverse environments with varying lighting conditions, textures, weather, and object categories. Therefore, accounting for and replanning with errors in maps is crucial when leveraging maps constructed with VLMs.
@@ 3 Method
In this section, we first introduce the background of our method in 3.1. Next, we illustrate the motivating concept behind our method to address a significant issue in current approaches in Section 3.2. We then detail our approach, which primarily focuses on leveraging uncertainty measures to select from a set of high-confidence candidates, as described in Section 3.3. Within this section, we further explain the single and multi-view uncertainty measures that our method utilizes. Subsequently, we outline the strategies we tested to generate the set of high-confidence candidates in Section 3.4. Finally, we describe the maps used in our experiments and discuss how they were adapted to integrate with our method in Section 3.5. Figure 2 illustrates a brief overview of our method.
@@
The conventional way to use navigation maps typically involves using a text query to retrieve the candidate with the highest matching score. Such a score is calculated with pretrained visual-language grounding models such as CLIP[8]. These grounding models have demonstrated impressive generalization abilities due to the large dataset used in the pertaining phase, which is probably why most of the works navigating with maps assume that the map is perfect. Nevertheless, we argue that despite the size of the training dataset, there might still be bias introduced by the model architecture or the dataset, which causes the user expectation of a query to diverge from the model’s belief. Furthermore, during the construction phase of the map, the model prediction might be interfered with by surrounding objects or some view-dependent bias, causing the extracted feature to be noisy.
@@ 3.3.1.1 Entropy:
Let $C=\{c_{1},c_{2},\ldots,c_{n}\}$ be the set of high-confidence candidates. For each $c_{i}$, we compute the entropy $H(c_{i})$ of its classification distribution $P_{i}=\{p_{1},p_{2},\ldots,p_{k}\}$ as:
@@
where $p_{j}$ is the probability of candidate $c_{i}$ belonging to class $j$. We select the candidate $c^{*}$ with the highest entropy:
@@
With this definition, a high entropy indicates that the classification distribution resembles a spread-out probability distribution, suggesting that the model considers several classes as plausible alternatives, thereby reflecting uncertainty. This approach ensures that we are not only considering candidates with high confidence but also those with high uncertainty, thereby reducing the likelihood of biased retrievals. For the specific method of selecting high-confidence candidates, please refer to Section 3.4.
@@
When the feature of a candidate is extracted by a small number of views, the resulting feature can easily be interfered by nearby objects or view-dependent model bias. To mitigate this issue, we propose to encourage multi-view consistency when retrieving the candidate. Similar to the single-view case, we first identify a list of high-confidence candidates and then encourage the retrieval of low-multi-view inconsistency (uncertainty) candidates. We experimented with two kinds of multi-view uncertainty measures: channel-average feature standard error and mean pairwise KL divergence on multi-view classification probabilities.
@@ 3.3.2.1 Channel-average Feature Standard Error:
where $\sigma_{k}$ is the standard deviation of the $k$-th feature across all views, $d$ is the dimensionality of the feature vector, and $m_{i}$ is the number of views for candidate $c_{i}$.
Finally, we choose the candidate $c^{*}$with the lowest standard error. Since most maps utilize the mean feature from multiple views for retrieval, a low standard error on these multi-view features indicates that the mean feature is statistically more reliable. Therefore encouraging low standard error could potentially reduce noisy retrievals.
@@ 3.3.2.2 Mean Pairwise KL Divergence:
We then choose the candidate $c^{*}$with the lowest mean pairwise KL divergence. A low mean pairwise KL divergence indicates that the classification distribution is consistent across multiple views. Similar to the standard error, promoting consistency in the classification distribution could help reduce noisy retrievals.
@@ 3.4.2 Top-k Category:
OpenMask3D[11] takes in the scene point cloud and posed RGB images, generates class-agnostic 3D masks on the point cloud, and uses the posed RGB images where the object is visible to provide semantic features. The features are calculated with CLIP with cropped images of the corresponding object and thus enable retrieval later on. If the object is visible in multiple views, OpenMask3D takes the average of multi-view features. The implementation details are further described in A.2.
@@
VLMaps takes the RGB-D image and its corresponding pose as input. It first generate the local point cloud which is then projected to the world coordinate frame and the map position with the information of camera poses. After that, the RGB image is fed into Visual Encoder of LSeg[10] to get its image feature, which will then be projected to the corresponding map position. Though we do not modify the process of building the map, we store some more data such as the standard error and KL divergence of each point on the map to decrease the size of the map and avoid repetitive computation in the evaluation process. When calculating the Channel-average Feature Standard Error, we incorporate the distance weighting [1, 5] information of each feature which is utilized when building the map. For more details, please refer to A.3.
@@ 4 Experiments
Table 1: OM3D Object Retrieval Success Rates. The upper half of the table presents the baseline methods, as described in 4.1.1, while the lower half displays variants of our method. The ’Replan Strategy’ column indicates the replanning strategies used, as detailed in 3.3. The ’Selection Criteria’ column specifies the criteria employed to generate the high-confidence candidate set, as described in 3.4. The best-performing entry for each column, excluding the Oracle, is bolded to emphasize the best-performing method given $k$. The best-performing entry for each row is highlighted to showcase the performance of each method when a hyper-parameter search on $k$ is available. Finally, the best-performing baseline entry is italicized for easier comparison.
@@
Table 2 showcases the effectiveness of various replanning strategies compared to the ”Max confidence (highest score)” baseline. The ”Max entropy” strategies demonstrate significant improvements over the baseline, reducing bias by considering entropy. The ”Min stderr” strategy consistently outperforms all other methods, including the baseline, by focusing on minimizing the standard error in confidence predictions. This approach proves to be the most robust and effective, highlighting the importance of incorporating uncertainty and multi-view consistency in replanning.
@@
Table 2: VLMaps Replanning Subgoal Success Rates. Similar to table 1, the baseline methods and ’Replan Strategy’ column are explained in 4.1.1 and 3.3, respectively. The ’Selection Criteria’ column describes the criteria for generating the high-confidence candidate set, as detailed in 3.4.
4.3 Ablation Studies: Different Choice on Choosing Maximal/Minimal Uncertainty.
To further support our assumptions, we conducted an ablation study with the OpenMask3D setting by examining different directions of the uncertainty metric. Results in figures 4 and 4 indicate that the variants consistently underperform the baselines.
Figure 3: Entropy vs. K. If a model is biased, selecting low-entropy candidates from it might further reinforce the bias and ultimately degrade performance. This figure illustrates the performance of choosing minimum entropy targets from top-k confidence candidates.
Figure 4: KL Divergence vs. K. When the multi-view prediction is noisy, it’s more likely to provide false information. This figure depicts the performance of selecting the least consistent candidate (i.e., maximum mean pairwise KL divergence) from top-k confidence candidates.
5 Limitation